Which expression is the factored form of 8x^2 - 22x + 5?
A. (4x - 5) (2x - 1)
B. (4x - 1) (2x - 5)
C. (4x + 1) (2x + 5)
D. (4x + 5) (2x + 1)

Answer by greenestamps(13200)



The constant term in the expression is positive, so the signs in both factors must be the same.

Given that positive constant term, since the sign of the linear term in the expression is negative, the signs in both factors must be negative. So we can eliminate answer choices C and D.

Answer choices A and B when multiplied together both give the correct "8x^2" and "+5" terms, so the correct answer is going to be the one that gives the correct middle term.

For answer choice A, the middle term of the product is (-5)(2)+(-1)(4) = (-10)(-4) = -14 -- not correct.

For answer choice B, the middle term of the product is (-1)(2)+(-5)(4) = (-2)+(-20) = -22 -- correct.

ANSWER: B